Find the probability of drawing a heart or a club from a standard deck of cards.: Find the probability of drawing a heart or a club from a standard deck of cards.
Answer by Fombitz(1740)

52 cards
13 clubs
13 hearts
26 clubs or hearts
P(club or heart)=26/52=1/2